# UniME-V2: MLLM-as-a-Judge for Universal Multimodal Embedding Learning

### v1

@@ Introduction
Multimodal embedding models aim to encode heterogeneous multimodal data into a unified dense representation space, enabling a wide range of downstream applications such as visual question answering (Dong et al. 2025; Hamza et al. 2025; Li et al. 2025) and multimodal retrieval (Zheng et al. 2025; Gu et al. 2025b; Yang et al. 2025; Gu et al. 2024). With the increasing adoption of these models, multimodal representation learning has garnered significant research attention. Among these models, CLIP (Radford et al. 2021) stands out as a pioneering approach, achieving remarkable performance in text-image retrieval by leveraging cross-modal contrastive learning on large-scale web-collected image-text pairs (Schuhmann et al. 2021). However, its effectiveness is hindered by three major limitations: (1) CLIP enforces a strict text token limit of 77, which restricts its ability to process detailed or lengthy descriptions (Zhang et al. 2024a; Cao, Wei, and Ma 2024; Huang et al. 2024b); (2) Its dual-encoder design processes images and text independently, which reduces its effectiveness in handling complex tasks, such as instruction-following multimodal retrieval (Jiang et al. 2025; Liu et al. 2024; Gu et al. 2025a); and (3) CLIP exhibits limited proficiency in advanced language understanding, struggles with handling compositionality, and often demonstrates bag-of-words behavior (Yuksekgonul et al. 2022; Tschannen et al. 2023; Hu et al. 2025).
@@
Recent advances in Large Language Models (LLMs) have achieved state-of-the-art performance on the MTEB benchmark (Muennighoff et al. 2022). Motivated by these developments (Lee et al. 2024; BehnamGhader et al. 2024), researchers are currently exploring how to utilize Multimodal Large Language Models (MLLMs) to learn universal multimodal representation. E5-V (Jiang et al. 2024) adopts a unimodal contrastive learning approach, training the language component of MLLMs on sentence pairs to better align the cross-modal representation spaces. VLM2Vec (Jiang et al. 2025) introduces the Massive Multimodal Embedding Benchmark (MMEB), comprising 36 datasets across four meta-tasks, and proposes a contrastive learning framework to repurpose pre-trained vision-language models as embedding models via training on the MMEB dataset. QQMM (Xue, Li, and Liu 2025a) provides an in-depth analysis of the gradients derived from the InfoNCE loss and proposes amplifying gradients associated with hard negative samples to encourage the model to learn more discriminative embeddings. UniME (Gu et al. 2025a) presents a two-stage framework that leverages a powerful LLM-based teacher model to improve the embedding capabilities of the language component in MLLMs. Furthermore, it incorporates a hard negative sampling strategy that selects multiple challenging negatives per instance within each batch. Despite these advances, existing methods fail to fully exploit the semantic differences between candidates and are limited by the lack of diversity in negative samples. Moreover, the raw embeddings produced by these models are frequently inadequate for reliably discriminating between hard negatives and false negatives.
@@
  • We introduce an MLLM-as-a-Judge pipeline for hard negative mining that uses the advanced understanding capabilities of MLLM to assess the semantic alignment of each query-candidate pair within a globally retrieved potential hard negative set.
@@
Multimodal Large Language Models (MLLMs) extend traditional LLMs to process and integrate information across multiple modalities (Wang et al. 2025b, a; Xie et al. 2024; Bai et al. 2025; An et al. 2025a, 2024). As a foundational contribution, LLaVA (Liu et al. 2023) leverages a subset of the CC3M (Changpinyo et al. 2021) dataset to achieve more balanced conceptual coverage. In this approach, the visual encoder and language model remain frozen, while only the projection layer is trained to align visual features with language tokens. Subsequently, numerous MLLM variants (Peng et al. 2023; Lin et al. 2024a; Tang et al. 2025a, b; An et al. 2025b) have achieved remarkable results in multimodal understanding and reasoning tasks. For instance, CogVLM (Wang et al. 2023) incorporates a trainable visual expert module into the attention and feed-forward layers of the language model, achieving significant improvements on 17 standard cross-modal benchmarks. Similarly, Qwen2-VL (Wang et al. 2024) introduces the Naive Dynamic Resolution mechanism and integrates M-RoPE to enhance positional information fusion, yielding competitive performance across diverse benchmarks. LLaVA-OneVision (Li et al. 2024) pushes the boundaries of open MLLMs by excelling in single-image, multi-image, and video tasks, showcasing robust video understanding through effective task transfer from image-based training. Although these advances have significantly improved the understanding capabilities of MLLMs, further research is needed to explore how MLLMs can effectively learn unified multimodal representations.
@@
CLIP (Radford et al. 2021) demonstrates strong image-text retrieval performance through large-scale cross-modal contrastive learning but faces three key limitations: (1) A 77-token text truncation restricts fine-grained semantic alignment (Zhang et al. 2024a; Cao, Wei, and Ma 2024; Huang et al. 2024b); (2) Its dual-encoder architecture limits effective cross-modal fusion, particularly for instruction-sensitive tasks (Jiang et al. 2025; Liu et al. 2024; Gu et al. 2025a); and (3) Simplistic language modeling results in bag-of-words representations (Yuksekgonul et al. 2022; Tschannen et al. 2023; Hu et al. 2025; Lei et al. 2024; Huang et al. 2024a; Andonian, Chen, and Hamid 2022). To address these issues, recent studies have incorporated MLLMs for enhanced multimodal representation learning. E5-V (Jiang et al. 2024) employs unimodal contrastive learning, training the language component of MLLMs on sentence pairs to reduce cross-modal representation gaps. VLM2Vec (Jiang et al. 2025) introduces the Massive Multimodal Embedding Benchmark (MMEB) and adapts state-of-the-art vision-language models into embedding models using a contrastive framework trained on MMEB. QQMM (Xue, Li, and Liu 2025a) analyzes InfoNCE loss gradients and proposes enhancing gradients associated with hard negatives to improve embedding discrimination. UniME (Gu et al. 2025a) adopts a two-stage framework with an LLM-based teacher model to refine language embeddings and employs a hard negative sampling strategy, selecting multiple challenging negatives per batch. Despite these advancements, existing methods still under-utilize the semantic differences among candidates and struggle to effectively identify and leverage hard negatives during retrieval.
@@
Previous works (Jiang et al. 2025; Liu et al. 2024) primarily rely on in-batch hard negative mining, where query-candidate embedding similarities are computed to sample negatives. However, this method often suffers from limited negative sample diversity and insufficient embedding discriminative power to effectively distinguish false and hard negatives. To overcome these challenges, as shown in Fig. 2, we first construct a potential hard negative set using global retrieval. After that, inspired by previous work (Zheng et al. 2023; Chen et al. 2024a), we leverage the robust understanding capabilities of MLLMs to assess the semantic alignment of each query-candidate pair and generate a soft semantic matching score. This score guides hard negative mining, enabling the identification of diverse and high-quality hard negatives while reducing the impact of false negatives.
@@
CLIP (ViT-L)(Jiang et al. 2025) 0.4B 42.8 9.1 53.0 51.8 37.1 38.7 39.2
OpenCLIP (ViT-L)(Radford et al. 2021) 0.4B 41.5 6.9 44.6 53.5 32.8 36.0 36.6
Magiclens (ViT-L)(Zhang et al. 2024b) 0.4B 38.8 8.3 35.4 26.0 31.0 23.7 27.1
SigLIP (So/14)(Zhai et al. 2023) 0.9B 40.3 8.4 31.6 59.5 32.3 38.0 35.0
BLIP2 (ViT-L)(Li et al. 2023) 1.2B 27.0 4.2 33.9 47.0 25.3 25.1 28.0
CLIP (ViT-BigG/14)(Cherti et al. 2022) 2.5B 52.3 14.0 50.5 60.3 38.9 45.8 44.3
EVA-CLIP(Sun et al. 2023) 8B 56.0 10.4 49.2 58.9 38.1 45.6 43.7
E5-V (Phi3.5-V)(Jiang et al. 2024) 4.2B 39.1 9.6 38.0 57.6 33.1 31.9 36.1
E5-V (LLaVA-1.6)(Jiang et al. 2024) 7B 39.7 10.8 39.4 60.2 34.2 33.4 37.5
@@
CLIP (ViT-L)(Jiang et al. 2025) 0.4B 55.2 19.7 53.2 62.2 47.6 42.8 47.6
VLM2Vec (Qwen2-VL)(Jiang et al. 2025) 2B 59.0 49.4 65.4 73.4 66.0 52.6 60.1
VLM2Vec (Qwen2-VL)(Jiang et al. 2025) 7B 62.6 57.8 69.9 81.7 72.2 57.8 65.8
LLaVE (LLaVA-OV)(Lan et al. 2025) 7B 65.7 65.4 70.9 91.9 75.0 64.4 70.3
QQMM (LLaVA-OV)(Xue, Li, and Liu 2025b) 7B 66.8 66.8 70.5 90.4 74.7 65.6 70.7
UniME (Qwen2-VL)(Gu et al. 2025a) 2B 59.0 53.4 64.9 69.6 65.5 54.6 60.6
UniME (Qwen2-VL)(Gu et al. 2025a) 7B 64.7 59.0 71.6 82.7 72.2 61.4 67.4
UniME (LLaVA-OV)(Gu et al. 2025a) 7B 66.8 66.6 70.5 90.9 74.6 65.8 70.7
@@
Table 1: Results on the MMEB benchmark. IND: in-distribution, OOD: out-of-distribution. Scores are average Precision@1. Detailed results in the supplementary material.
@@
UniME-V2. Previous works (Jiang et al. 2025; Gu et al. 2025a) are limited by a rigid one-to-one mapping, which restricts the ability to learn distinctions among diverse negative samples. To address this, as shown in Fig. 3, we propose an MLLM judgment based distribution alignment framework, leveraging soft semantic matching scores as supervised signals to improve representation performance. Specifically, given a query $q$ and its candidate set $\Omega_{c}=\{c_{t},c_{1},...,c_{k}\}$, we input them into the MLLM and extract the last token as embeddings for the query $e_{q}$ and candidates $\mathrm{E_{c}}=\{e_{c}^{+},e_{c_{1}}^{-},..,e_{c_{k}}^{-}\}$, where $e_{c}^{+}$ is the embedding of target candidate and $k$ is the hard negative number for each query. We then compute the relation score matrix between the query embedding $e_{q}$ and candidate embeddings $\mathrm{E_{c}}$ as follows:
@@
To enhance learning robustness and ensure matrix symmetry, we employ JS-Divergence, a symmetric alternative to KL-Divergence (Nielsen 2020). The final loss function $\mathcal{L}$ is defined as:
@@
UniME-V2-Reranker. Following previous works (Liu et al. 2024; Lin et al. 2024b), we train a reranking model to enhance retrieval precision following initial embedding-based retrieval. Specifically, we train UniME-V2-Reranker using joint pairwise and listwise approaches to enhance its reranking capability (refer to Fig. 3). In pairwise training, we construct two pairs for each query $q$ by combining with the positive candidate $c_{t}$ and the hardest negatives $c_{h}$. We then instruct UniME-V2-Reranker to output YES for the positive and NO for the negative. The pairwise loss $\mathcal{L}_{pair}$ is computed using the cross-entropy loss function as:
@@
Following VLM2Vec (Jiang et al. 2025) and UniME (Gu et al. 2025a), we employ 20 in-distribution datasets from the MMEB benchmark, which cover four core multimodal tasks: classification, visual question answering, multimodal retrieval, and visual grounding. This comprehensive training corpus, incorporating both unimodal and multimodal input data, totals 662k carefully curated training pairs, ensuring robust model adaptation across diverse multimodal tasks.
@@
OpenCLIP (ViT-L) (Radford et al. 2021) 0.4B 67.3 87.2 37.0 58.1 81.8 84.0 47.0 47.0 79.5 62.7 74.9
CLIP (ViT-BigG/14) (Cherti et al. 2022) 2.5B 79.5 92.9 51.3 67.3 90.1 93.6 77.8 80.7 86.5 68.9 88.4
EVA-CLIP (Sun et al. 2023) 8B 80.3 94.5 52.0 70.1 93.1 91.2 80.4 77.8 85.9 70.3 86.7
E5-V (Phi3.5-V) (Jiang et al. 2024) 4.2B 72.2 79.6 44.7 53.4 86.0 88.5 83.8 83.6 88.2 66.6 75.3
E5-V (LLaVA-1.6) (Jiang et al. 2024) 7B 77.3 85.7 49.1 57.6 85.1 82.1 88.9 83.2 86.3 68.7 66.9
VLM2Vec (Qwen2-VL) (Jiang et al. 2025) 2B 69.3 89.6 40.0 62.5 78.1 88.2 78.7 83.9 67.2 46.5 66.4
VLM2Vec (Qwen2-VL) (Jiang et al. 2025) 7B 80.0 94.2 49.2 68.5 78.5 90.4 94.0 94.2 70.0 51.7 72.2
UniME (Qwen2-VL) (Gu et al. 2025a) 2B 74.9 90.6 44.0 63.5 83.6 88.6 83.3 83.2 65.6 45.2 65.7
UniME (Qwen2-VL) (Gu et al. 2025a) 7B 80.8 92.7 50.9 69.8 86.5 93.8 95.3 94.0 68.8 53.0 69.8
UniME (LLaVA-OV) (Gu et al. 2025a) 7B 83.3 94.4 54.8 74.0 93.9 89.3 94.3 95.5 80.5 65.5 82.2
@@
In this study, we evaluate UniME-V2 across both in-distribution (20 test sets) and out-of-distribution (16 test sets) benchmarks from MMEB (Jiang et al. 2025) to assess its multimodal embedding capabilities across diverse retrieval tasks. Following standard evaluation protocols (Liu et al. 2024; Jiang et al. 2025), we report Precision, measuring the proportion of correct matches among the top-ranked candidates for each dataset. To further examine the unimodal embedding performance of UniME-V2, we conduct experiments on multiple cross-modal retrieval tasks, including short-caption image-text retrieval on Flickr30K (Plummer et al. 2015) and COCO2014 (Lin et al. 2014), long-caption image-text retrieval on ShareGPT4V (Chen et al. 2024b) and Urban1K (Zhang et al. 2024a), and compositional retrieval on SugarCrepe (Hsieh et al. 2023). Consistent with the MMEB benchmark, we use Precision as the primary evaluation metric across all datasets.

### v2

@@ Introduction
Multimodal embedding models aim to encode heterogeneous multimodal data into a unified dense representation space, enabling a wide range of downstream applications such as visual question answering (dong2024toward; hamza2025llava; li2025miv) and multimodal retrieval (zheng2025gradient; gu2025realsyn; yang2025clip; gu2024rwkv). With the increasing adoption of these models, multimodal representation learning has garnered significant research attention. Among these models, CLIP (CLIP) stands out as a pioneering approach, achieving remarkable performance in text-image retrieval by leveraging cross-modal contrastive learning on large-scale web-collected image-text pairs (schuhmann2021laion). However, its effectiveness is hindered by three major limitations: (1) CLIP enforces a strict text token limit of 77, which restricts its ability to process detailed or lengthy descriptions (longclip; flame; llm2clip); (2) Its dual-encoder design processes images and text independently, which reduces its effectiveness in handling complex tasks, such as instruction-following multimodal retrieval (VLM2Vec; LamRA; unime); and (3) CLIP exhibits limited proficiency in advanced language understanding, struggles with handling compositionality, and often demonstrates bag-of-words behavior (negclip; tschannen2023image; degla).
@@
Recent advances in Large Language Models (LLMs) have achieved state-of-the-art performance on the MTEB benchmark (mteb). Motivated by these developments (nvembed; llm2vec), researchers are currently exploring how to utilize Multimodal Large Language Models (MLLMs) to learn universal multimodal representation. E5-V (E5V) adopts a unimodal contrastive learning approach, training the language component of MLLMs on sentence pairs to better align the cross-modal representation spaces. VLM2Vec (VLM2Vec) introduces the Massive Multimodal Embedding Benchmark (MMEB), comprising 36 datasets across four meta-tasks, and proposes a contrastive learning framework to repurpose pre-trained vision-language models as embedding models via training on the MMEB dataset. QQMM (xue2025improve) provides an in-depth analysis of the gradients derived from the InfoNCE loss and proposes amplifying gradients associated with hard negative samples to encourage the model to learn more discriminative embeddings. UniME (unime) presents a two-stage framework that leverages a powerful LLM-based teacher model to improve the embedding capabilities of the language component in MLLMs. Furthermore, it incorporates a hard negative sampling strategy that selects multiple challenging negatives per instance within each batch. Despite these advances, existing methods fail to fully exploit the semantic differences between candidates and are limited by the lack of diversity in negative samples. Moreover, the raw embeddings produced by these models are frequently inadequate for reliably discriminating between hard negatives and false negatives.
@@
  • We introduce a MLLM-as-a-Judge pipeline for hard negative mining that uses the advanced understanding capabilities of MLLM to assess the semantic alignment of each query-candidate pair within a globally retrieved potential hard negative set.
@@
Multimodal Large Language Models (MLLMs) extend traditional LLMs to process and integrate information across multiple modalities (wang2025vl; wang2025code; xie2024croc; bai2025qwen2; an2025unictokens; an2024mc). As a foundational contribution, LLaVA (llava) leverages a subset of the CC3M (cc3m) dataset to achieve more balanced conceptual coverage. In this approach, the visual encoder and language model remain frozen, while only the projection layer is trained to align visual features with language tokens. Subsequently, numerous MLLM variants (kosmos; vila; tang2025intervening; tang2025seeing; an2025llava) have achieved remarkable results in multimodal understanding and reasoning tasks. For instance, CogVLM (wang2023cogvlm) incorporates a trainable visual expert module into the attention and feed-forward layers of the language model, achieving significant improvements on 17 standard cross-modal benchmarks. Similarly, Qwen2-VL (wang2024qwen2) introduces the Naive Dynamic Resolution mechanism and integrates M-RoPE to enhance positional information fusion, yielding competitive performance across diverse benchmarks. LLaVA-OneVision (li2024llava) pushes the boundaries of open MLLMs by excelling in single-image, multi-image, and video tasks, showcasing robust video understanding through effective task transfer from image-based training. Although these advances have significantly improved the understanding capabilities of MLLMs, further research is needed to explore how MLLMs can effectively learn unified multimodal representations.
@@
CLIP (CLIP) demonstrates strong image-text retrieval performance through large-scale cross-modal contrastive learning but faces three key limitations: (1) A 77-token text truncation restricts fine-grained semantic alignment (longclip; flame; llm2clip); (2) Its dual-encoder architecture limits effective cross-modal fusion, particularly for instruction-sensitive tasks (VLM2Vec; LamRA; unime); and (3) Simplistic language modeling results in bag-of-words representations (negclip; tschannen2023image; degla; lei2023mcad; huang2024cross; andonian2022robust). To address these issues, recent studies have incorporated MLLMs for enhanced multimodal representation learning. E5-V (E5V) employs unimodal contrastive learning, training the language component of MLLMs on sentence pairs to reduce cross-modal representation gaps. VLM2Vec (VLM2Vec) introduces the Massive Multimodal Embedding Benchmark (MMEB) and adapts state-of-the-art vision-language models into embedding models using a contrastive framework trained on MMEB. QQMM (xue2025improve) analyzes InfoNCE loss gradients and proposes enhancing gradients associated with hard negatives to improve embedding discrimination. UniME (unime) adopts a two-stage framework with an LLM-based teacher model to refine language embeddings and employs a hard negative sampling strategy, selecting multiple challenging negatives per batch. Despite these advancements, existing methods still under-utilize the semantic differences among candidates and struggle to effectively identify and leverage hard negatives during retrieval.
@@
Previous works (VLM2Vec; LamRA) primarily rely on in-batch hard negative mining, where query-candidate embedding similarities are computed to sample negatives. However, this method often suffers from limited negative sample diversity and insufficient embedding discriminative power to effectively distinguish false and hard negatives. To overcome these challenges, as shown in Fig. 2, we first construct a potential hard negative set using global retrieval. After that, inspired by previous work (zheng2023judging; chen2024mllm), we leverage the robust understanding capabilities of MLLMs to assess the semantic alignment of each query-candidate pair and generate a soft semantic matching score. This score guides hard negative mining, enabling the identification of diverse and high-quality hard negatives while reducing the impact of false negatives.
@@
CLIP (ViT-L)(VLM2Vec) 0.4B 42.8 9.1 53.0 51.8 37.1 38.7 39.2
OpenCLIP (ViT-L)(CLIP) 0.4B 41.5 6.9 44.6 53.5 32.8 36.0 36.6
Magiclens (ViT-L)(magiclens) 0.4B 38.8 8.3 35.4 26.0 31.0 23.7 27.1
SigLIP (So/14)(siglip) 0.9B 40.3 8.4 31.6 59.5 32.3 38.0 35.0
BLIP2 (ViT-L)(blip2) 1.2B 27.0 4.2 33.9 47.0 25.3 25.1 28.0
CLIP (ViT-BigG/14)(CLIP_bigG14) 2.5B 52.3 14.0 50.5 60.3 38.9 45.8 44.3
EVA-CLIP(EVA_CLIP_18B) 8B 56.0 10.4 49.2 58.9 38.1 45.6 43.7
E5-V (Phi3.5-V)(E5V) 4.2B 39.1 9.6 38.0 57.6 33.1 31.9 36.1
E5-V (LLaVA-1.6)(E5V) 7B 39.7 10.8 39.4 60.2 34.2 33.4 37.5
@@
CLIP (ViT-L)(VLM2Vec) 0.4B 55.2 19.7 53.2 62.2 47.6 42.8 47.6
VLM2Vec (Qwen2-VL)(VLM2Vec) 2B 59.0 49.4 65.4 73.4 66.0 52.6 60.1
VLM2Vec (Qwen2-VL)(VLM2Vec) 7B 62.6 57.8 69.9 81.7 72.2 57.8 65.8
LLaVE (LLaVA-OV)(lan2025llave) 7B 65.7 65.4 70.9 91.9 75.0 64.4 70.3
QQMM (LLaVA-OV)(qqmm) 7B 66.8 66.8 70.5 90.4 74.7 65.6 70.7
UniME (Qwen2-VL)(unime) 2B 59.0 53.4 64.9 69.6 65.5 54.6 60.6
UniME (Qwen2-VL)(unime) 7B 64.7 59.0 71.6 82.7 72.2 61.4 67.4
UniME (LLaVA-OV)(unime) 7B 66.8 66.6 70.5 90.9 74.6 65.8 70.7
@@
Table 1: Results on the MMEB benchmark. IND: in-distribution, OOD: out-of-distribution. Scores are average Precision@1. Detailed results are in the supplementary material.
@@
UniME-V2. Previous works (VLM2Vec; unime) are limited by a rigid one-to-one mapping, which restricts the ability to learn distinctions among diverse negative samples. To address this, as shown in Fig. 3, we propose an MLLM judgment based distribution alignment framework, leveraging soft semantic matching scores as supervised signals to improve representation performance. Specifically, given a query $q$ and its candidate set $\Omega_{c}=\{c_{t},c_{1},...,c_{k}\}$, we input them into the MLLM and extract the last token as embeddings for the query $e_{q}$ and candidates $\mathrm{E_{c}}=\{e_{c}^{+},e_{c_{1}}^{-},..,e_{c_{k}}^{-}\}$, where $e_{c}^{+}$ is the embedding of target candidate and $k$ is the hard negative number for each query. We then compute the relation score matrix between the query embedding $e_{q}$ and candidate embeddings $\mathrm{E_{c}}$ as follows:
@@
To enhance learning robustness and ensure matrix symmetry, we employ JS-Divergence, a symmetric alternative to KL-Divergence (nielsen2020generalization). The final loss function $\mathcal{L}$ is defined as:
@@
UniME-V2-Reranker. Following previous works (LamRA; mm_embed), we train a reranking model to enhance retrieval precision following initial embedding-based retrieval. Specifically, we train UniME-V2-Reranker using joint pairwise and listwise approaches to enhance its reranking capability (refer to Fig. 3). In pairwise training, we construct two pairs for each query $q$ by combining with the positive candidate $c_{t}$ and the hardest negatives $c_{h}$. We then instruct UniME-V2-Reranker to output YES for the positive and NO for the negative. The pairwise loss $\mathcal{L}_{pair}$ is computed using the cross-entropy loss function as:
@@
Following VLM2Vec (VLM2Vec) and UniME (unime), we employ 20 in-distribution datasets from the MMEB benchmark, which cover four core multimodal tasks: classification, visual question answering, multimodal retrieval, and visual grounding. This comprehensive training corpus, incorporating both unimodal and multimodal input data, totals 662k carefully curated training pairs, ensuring robust model adaptation across diverse multimodal tasks.
@@
OpenCLIP (ViT-L) (CLIP) 0.4B 67.3 87.2 37.0 58.1 81.8 84.0 47.0 47.0 79.5 62.7 74.9
CLIP (ViT-BigG/14) (CLIP_bigG14) 2.5B 79.5 92.9 51.3 67.3 90.1 93.6 77.8 80.7 86.5 68.9 88.4
EVA-CLIP (EVA_CLIP_18B) 8B 80.3 94.5 52.0 70.1 93.1 91.2 80.4 77.8 85.9 70.3 86.7
E5-V (Phi3.5-V) (E5V) 4.2B 72.2 79.6 44.7 53.4 86.0 88.5 83.8 83.6 88.2 66.6 75.3
E5-V (LLaVA-1.6) (E5V) 7B 77.3 85.7 49.1 57.6 85.1 82.1 88.9 83.2 86.3 68.7 66.9
VLM2Vec (Qwen2-VL) (VLM2Vec) 2B 69.3 89.6 40.0 62.5 78.1 88.2 78.7 83.9 67.2 46.5 66.4
VLM2Vec (Qwen2-VL) (VLM2Vec) 7B 80.0 94.2 49.2 68.5 78.5 90.4 94.0 94.2 70.0 51.7 72.2
UniME (Qwen2-VL) (unime) 2B 74.9 90.6 44.0 63.5 83.6 88.6 83.3 83.2 65.6 45.2 65.7
UniME (Qwen2-VL) (unime) 7B 80.8 92.7 50.9 69.8 86.5 93.8 95.3 94.0 68.8 53.0 69.8
UniME (LLaVA-OV) (unime) 7B 83.3 94.4 54.8 74.0 93.9 89.3 94.3 95.5 80.5 65.5 82.2
@@
In this study, we evaluate UniME-V2 across both in-distribution (20 test sets) and out-of-distribution (16 test sets) benchmarks from MMEB (VLM2Vec) to assess its multimodal embedding capabilities across diverse retrieval tasks. Following standard evaluation protocols (LamRA; VLM2Vec), we report Precision, measuring the proportion of correct matches among the top-ranked candidates for each dataset. To further examine the unimodal embedding performance of UniME-V2, we conduct experiments on multiple cross-modal retrieval tasks, including short-caption image-text retrieval on Flickr30K (flickr30k) and COCO2014 (MSCOCO2014), long-caption image-text retrieval on ShareGPT4V (sharegpt4v) and Urban1K (longclip), and compositional retrieval on SugarCrepe (sugarcrepe). Consistent with the MMEB benchmark, we use Precision as the primary evaluation metric across all datasets.